# Sampling-based Model Predictive Control Leveraging Parallelizable Physics Simulations

### v2

@@ III-C2 Extension to different robots
Box
Pose A
5
9.66 $\pm$ 0.84
Pose B
5
12.84 $\pm$ 0.564
Sphere
Pose A
5
8.76 $\pm$ 0.38
Pose B
5
7.45 $\pm$ 0.59
@@
(b) Pushing to the goal on the left with ${}^{\circ}$ rotation.
Figure 10: Pushing to two different goals with a 7 DOF manipulator directly controlling all joint velocities. Our method allows the re-positioning of the end-effector around the object without specifying any particular desired contact point.
@@
We perform differential drive non-prehensile pushing, see Fig. 9, with the same cost function as before (see 11) but re-tuned. One can indeed change the robot for the task by simply changing the URDF, neglecting all the additional contact modeling that would be required in a classical model-based MPC. The time taken to push the box to the goal was 18.31s. In the mobile non-prehensile pushing experiments, objects to manipulate are spawned with 30% uncertainty on mass and friction sampled uniformly, while object size is randomized with Gaussian noise with a standard deviation of 5mm.
@@ III-D Real-world experiments
To demonstrate the applicability of our approach, we transfer to the real world a subset of the non-prehensile manipulation tasks previously presented in Section III-C with both the robot manipulator and the omnidirectional base. In particular, in Fig. 10, we show the results of the 7 DOF manipulator pushing a product to two different goals, similar to the simulations corresponding to Table II. As presented in Fig. 1, the samples are rolled out in $K=500$ simulated environments in IsaacGym which, at each timestep, are initialized to the state of the real world. Based on this, the optimal control is estimated and applied to the real system.
When transferring to the real world, compared to the experiments in Section III-C1, only the weights of the cost function were re-tuned. The horizon, control frequency, number of samples, structure of the cost function, and randomization of the sampled environments remained unchanged.
From the experimental evaluation on the real robot, we observe that the time to complete the different pushing tasks and the final position error are comparable to the results in the simulation from Table II. Importantly, these results are achieved without taking assumptions on specific contact points, thus the robot can naturally re-position itself and change contact location autonomously. Additionally, our method allows to sample joint velocities directly thus we do not restrict the sampling to 2D end-effector trajectories to be then translated into joint commands as often seen in other approaches.
Lastly, to demonstrate robustness, we disturb by hand the execution of pushing tasks with the manipulator and the omnidirectional base. Since we do not assume the robot to be behind the object to be pushed for successful execution, and since the planning and execution happen in real-time at $25Hz$, we can largely perturb the task and let the robot compensate.
@@ IV Discussion
In this section, we discuss key aspects and potential future work related to our solution. First, the computational demands of planning and control with our method can be high when extending the time horizon to several seconds. To keep the time horizon limited for real-time control while preventing being trapped in local minima, future work should incorporate global planning techniques such as A*, RRT, and Probabilistic Roadmaps (PRM) [26] to guide the local planner. Similarly to warm starting predictive controllers [5, 6], one could make use of motion libraries of previous executions or learned policies along with random rollouts, to improve the sampling efficiency and exploration. Second, in real-world scenarios, uncertainties and discrepancies between simulated and actual environments could present challenges for achieving precise movements and manipulation. We utilized randomization of object properties in the rollouts to address some uncertainties. However, online system identification to converge to the true model parameters is not performed. Enhancing the robustness of the MPPI algorithm itself by reducing model uncertainty, as demonstrated by [18], could further improve performance. Third, tuning control algorithms for optimal performance is time-consuming. However, implementing autotuning techniques can automate the process and reduce manual effort. Finally, incorporating additional sensor support such as lidars and signed distance fields could be highly beneficial. This would enable, for instance, collision avoidance without the need for complex simulated worlds.

### v3

@@ III-C2 Extension to different robots
Box Pose A 5 9.66 $\pm$ 0.84
Pose B 5 12.84 $\pm$ 0.564
Sphere Pose A 5 8.76 $\pm$ 0.38
Pose B 5 7.45 $\pm$ 0.59
@@
(b) Pushing to the goal on the left with 90 rotation.
Figure 10: Pushing to two goals with a 7 DOF manipulator directly controlling all joint velocities. Our method allows the end-effector to re-position around the object without specifying any desired contact point.
@@
We perform differential drive non-prehensile pushing, see Fig. 9, with the same cost function as before (see 11) but re-tuned. One can change the robot for the task by changing the URDF, neglecting all the additional contact modeling required in a classical model-based MPC. The time taken to push the box to the goal was 18.31s. In the mobile non-prehensile pushing experiments, objects to manipulate are spawned with 30% uncertainty on mass and friction sampled uniformly, while object size is randomized with Gaussian noise with a standard deviation of 5mm.
@@ III-D Real-world experiments
To demonstrate the applicability of our approach, we transfer to the real world a subset of the non-prehensile manipulation tasks previously presented in Section III-C with both the robot manipulator and the omnidirectional base. In particular, in Fig. 10, we show the results of the 7 DOF manipulator pushing a product to two different goals, similar to the simulations corresponding to Table II. As presented in Fig. 1, the samples are rolled out in $K=500$ simulated environments in IsaacGym, which, at each timestep, are initialized to the state of the real world. Based on this, the optimal control is estimated and applied to the real system.
When transferring to the real world, compared to the experiments in Section III-C1, only the cost function weights were re-tuned. The horizon, control frequency, number of samples, structure of the cost function, and randomization of the sampled environments remained unchanged.
From the experimental evaluation on the real robot, we observe that the time to complete the pushing tasks and the final position errors are comparable to the results in the simulation from Table II. Importantly, these results are achieved without making assumptions on specific contact points. Thus, the robot can naturally re-position itself and change contact location autonomously. Additionally, our method allows us to sample joint velocities directly; thus, we do not restrict the sampling to 2D end-effector trajectories to be translated into joint commands, as often seen in other approaches.
Lastly, to demonstrate robustness, we disturb the execution of pushing tasks by hand with the manipulator and the omnidirectional base (Fig. 11). Since we do not assume the robot to be behind the object to be pushed for successful execution, and since the planning and execution happen in real-time at $25Hz$, we can largely perturb the task and let the robot compensate.
@@ IV Discussion
In this section, we discuss key aspects and potential future work related to our solution. First, the computational demands of planning and control with our method can be high when extending the time horizon to several seconds. To keep the time horizon limited for real-time control while preventing being trapped in local minima, future work should incorporate global planning techniques such as A*, RRT, and Probabilistic Roadmaps (PRM) [27] to guide the local planner. Similarly to warm starting predictive controllers [5, 6], one could make use of motion libraries of previous executions or learned policies along with random rollouts, to improve the sampling efficiency and exploration [28]. Second, in real-world scenarios, uncertainties and discrepancies between simulated and actual environments could present challenges for achieving precise movements and manipulation. We utilized randomization of object properties in the rollouts to address some uncertainties. However, online system identification to converge to the true model parameters is not performed. Enhancing the robustness of the MPPI algorithm itself by reducing model uncertainty, as demonstrated by [18], could further improve performance. Third, tuning control algorithms for optimal performance is time-consuming. Implementing autotuning techniques can automate the process and reduce manual effort. Finally, incorporating additional sensor support, such as lidars and signed distance fields, could be beneficial.